# ZMP-HH-25/5

### v1

@@ 4.3.2 $\text{rk}(\Lambda_{\rm trans})=3$ or $2$ ($\rho=18$ or $19$)
$\mathcal{M}=T^{2}_{U=T}\times\left(T^{2}_{(\mathbf{b}^{0},\mathbf{b}^{3})}\to\mathbb{C}_{\mathbf{z}_{0}}\right)\times\mathbb{C}^{*}_{u=\Phi e^{\widetilde{\mathbf{b}}^{0}}}\,,$ (4.40)
@@
$\mathcal{M}=T^{2}_{T=U=\tau_{0}}\times\left(T^{2}_{(\mathbf{b}^{0},\mathbf{b}^{3})}\to\mathbb{C}_{\mathbf{z}_{0}}\right)\times\mathbb{C}^{*}_{u=\Phi e^{{\widetilde{\mathbf{b}}^{0}}}}\,,$ (4.41)
where the complex structure and Kähler parameter of the torus are fixed to either $\tau_{0}=i$ or $\tau_{0}=(-1)^{1/3}$. For $\text{rk}(\Lambda_{\rm trans})=2$ we hence obtain a $\sigma$-model corresponding to the heterotic string on a local K3 times a torus with Kähler and complex structure fixed to one of the orbifold points in the fundamental domain of $\mathrm{SL}(2,\mathbb{Z})$.
@@ 4.3.3 $\text{rk}(\Lambda_{\rm trans})=21$ ($\rho=0$)
The final case to discuss is $\rho=0$ (i.e. $b=19$), corresponding to $\text{rk}(\Lambda_{\rm trans})=21$ and $\text{rk}(\Lambda_{\rm pol})=1$. The real scalar field $\mathbf{b}^{3}$ defined in (4.36) now has a free left-moving and an interacting right-moving piece. This leads to a left-right asymmetric worldsheet theory $\mathbf{b}^{3}$, which cannot correspond to the coordinate on a geometric $T^{2}$-fibre of a K3 surface. Accordingly, in $\mathcal{M}$ the geometric $(T^{2}_{(\mathbf{b}^{0},\mathbf{b}^{3})}\to\mathbb{C}_{\mathbf{z}_{0}})$-factor has to be replaced by a non-geometric CFT. Since $b_{-}^{3}$ is now a free field, this CFT must realise an additional left-moving ${\rm U}(1)$ current algebra on the worldsheet. In the bulk, the total gauge group is therefore a maximal rank subgroup of $E_{8}\times E_{8}\times\mathrm{U}(1)^{5}$ (or of $\mathrm{Spin}(32)/\mathbb{Z}_{2}\times\mathrm{U}(1)^{5}$). The extra ${\rm U}(1)$ factor suggests that the dynamics is captured by a CFT whose target is at an orbifold point, e.g. by fixing the complex structure of the torus fibre of $T^{2}_{(\mathbf{b}^{0},\mathbf{b}^{3})}\to\mathbb{C}_{\mathbf{z}_{0}}$ to $\tau=i$ which is constant over $\mathbb{C}_{\mathbf{z}_{0}}$ such that the monodromy for the $T^{2}$-fibre only generates a $\mathbb{Z}_{4}$ sub-group of $\mathrm{SL}(2,\mathbb{Z})$.
@@
The K3 surface that appears in the target space of the $\sigma$-model on the string worldsheet is non-compact since the base of its elliptic fibration is the complex plane $\mathbb{C}_{\mathbf{z}_{0}}$. On the dual heterotic side one might have expected instead a compactification on a compact K3 that is an elliptic fibration over $\mathbb{P}^{1}$ as opposed to the complex plane $\mathbb{C}_{\mathbf{z}_{0}}$. However, our analysis applies to the string at a point in the moduli space corresponding to a type II limit realised as a Tyurin degeneration. This maps to the tensionless limit for the string in Planck units, cf. (3.29). In the dual heterotic frame, the tensionless limit amounts to taking
@@
Thus in Planck units the volume of all curves on K3het diverges in the tensionless limit for the heterotic string, including in particular the base $\mathbb{P}^{1}$ of the elliptically fibred K3het. This explains why the analysis of the $\sigma$-model in the Tyurin degeneration only detects a non-compact base of the target space K3het. By the same reasoning, the volume of the torus fibre of K3het in Planck units must diverge in the emergent string limit. This perfectly matches the divergence of the moduli space metric, (4.19), for the interacting zero modes of the $\sigma$-model, including the zero modes $(\mathbf{b}^{0},\mathbf{b}^{3})$ associated with the elliptic fibre.
@@ 4.3.5 BPS invariants and modularity
Having established that the EFT string localised on the divisor $Z$ of a Tyurin degeneration is a critical heterotic string, we come back to the interpretation of the BPS invariants associated with special Lagrangian 3-cycles dual to elements in $H^{3}(V,\mathbb{Z})\cap{\rm Gr}_{2}(\Delta)$. As explained in Section 3.1, these are the 3-cycles of asymptotically vanishing volume in the infinite distance limit. To each such special Lagrangian 3-cycle $\Gamma_{0}$ one can associate a 2-cycle $C_{0}$ on $Z$ whose class lies in the transcendental lattice $\Lambda_{\rm trans}$ of the K3 surface $Z$. The heterotic string interpretation now suggests that the BPS indices $\Omega_{\rm BPS}(\Gamma_{0})$ count BPS states associated with the winding and Kaluza-Klein states of the heterotic string with respect to the torus factors on the dual heterotic target space ${\cal M}$. To see this, recall that in a type IIb degeneration, $\Lambda_{\rm trans}$ is a sublattice of signature $(2,b)$ of the K3 lattice $\Gamma_{(3,19)}$,
@@
where $U$ is the hyperbolic lattice of signature $(1,1)$ and $\Gamma_{16}$ collectively refers to the $E_{8}\times E_{8}$ or $\mathrm{Spin}(32)/\mathbb{Z}_{2}$ lattice. For example, for $b\geq 2$, $\Lambda_{\rm trans}$ can be embedded into a lattice containing two copies of $U$. The projection of the curve $C_{0}$ to each such hyperbolic factor then counts the wrapping and winding number of a heterotic string state with respect to the two one-cycles in the torus factor of the heterotic target space ${\cal M}$, while the projection on the $\Gamma_{16}$ factor accounts for the perturbative $E_{8}\times E_{8}$ or $\mathrm{Spin(32)}/\mathbb{Z}_{2}$ charge. This generalises also to $b=1$ or $b=0$: The only difference is that the momentum and winding along one or both of the torus 1-cycles are not independent, along the lines explained in Section 4.3.2. As first observed in [80] by computing the threshold corrections induced by the running of these states in the supergravity, the BPS indices of the winding or KK states are counted by a generating function with certain modular properties. In absence of space-time filling NS5-branes in the heterotic target space, the generating function is a meromorphic modular form. In the presence of NS5-branes, experience from similar setups, e.g. [81, 82, 83], suggests the appearance of modular anomalies which reflect incomplete cancellations in the BPS index. These are due to the backreaction of the NS5-branes on the heterotic $\sigma$-models. In the worst case this is expected to give rise to mock modular forms, which includes the quasi-modular forms [84, 85] that appear for the elliptic genus of the 5d/6d heterotic string, see [9, 86, 87] and references therein.
@@
with ${\cal I}$ a suitable subset of $\mathbb{Q}$, whose coefficients count the 4d $\mathcal{N}=2$ BPS invariants associated with $\Gamma_{0}$ according to the identification
@@
where $C_{0}=\partial_{*}\iota_{*}\Gamma_{0}\in\Lambda_{\rm trans}$ denotes the curve on $Z$ over which the asymptotic form of $\Gamma_{0}$ is fibered close to the Tyurin degeneration (cf. the discussion around (3.9)). If ${\rm Gr}_{3}=\emptyset$, $\theta(q)$ is in fact modular.
@@ 5.1 Worldsheet Theory of NS5-Brane String in Type IIA
Let us consider Type IIA string theory on a Calabi–Yau threefold $\widehat{V}$ that admits a K3-fibration over ^$\widehat{\mathbb{P}}^{1}_{b}$ with generic K3-fibre class $[\widehat{Z}]\in H_{4}(\widehat{V})$,
$\widehat{Z}\ \stackrel{{\scriptstyle\widehat{\iota}}}{{\hookrightarrow}}$ $\ \ \widehat{V}$ (5.1)
@@
^$\ \ \widehat{\mathbb{P}}^{1}_{b}\,.$
The limit in the Kähler moduli space in which ^$\widehat{\mathbb{P}}^{1}_{b}$ becomes large while the volume of the generic fibre remains constant corresponds to an emergent string limit. The emergent string in this case is given by the NS5-brane wrapping the generic K3-fibre and extended along two of the non-compact directions [8]. The situation is depicted in Figure -956.
Figure -956: A geometric depiction of the set-up described around equation (5.1). The Type IIA NS5-brane (indicated in green) is located at a generic point in the base ^$\widehat{\mathbb{P}}^{1}_{b}$ and wraps the generic K3-fibre $\widehat{Z}$. The emergent string is obtained in the limit where the volume of the base becomes large, while the volume of the generic fibre stays constant.
@@
For the string obtained by wrapping the Type IIA NS5-brane on $\widehat{Z}$, the massless spectrum on the worldsheet is obtained by dimensional reduction. The space transverse to $\widehat{Z}$ is topologically ^$\widehat{\mathbb{P}}^{1}_{b}\times\mathbb{C}$, where ^$\widehat{\mathbb{P}}^{1}_{b}$ is the base of the K3-fibration for $\widehat{V}$. Thus, two of the non-compact, real scalars on the worldvolume theory of the NS5-branes become compact and combine into the complex coordinate $\widehat{\mathbf{z}}\in\widehat{\mathbb{P}}^{1}_{b}$. The other two real scalars form a complex coordinate $\widehat{u}\in\mathbb{C}$. In addition, the compact scalar on the worldvolume theory of the NS5-branes remains in the spectrum after dimensional reduction. In the following, we refer to this scalar as $\widehat{\mathbf{b}}^{1}$. On the string worldsheet, these scalar fields split into left- and right-moving scalars, with the latter having fermionic superpartners to form multiplets of the $\mathcal{N}=(0,4)$ super-algebra on the worldsheet. In addition to the scalar modes on the 2d worldsheet arising from the scalar modes already present in the 6d $\mathcal{N}=(0,2)$ parent theory, the self-dual 2-form $\widehat{B}_{2}$ yields further scalar modes when expanded in 2-forms on $\widehat{Z}$. These scalars are compact. As for the MSW-string discussed in [88], $\widehat{B}_{2}$ gives rise to $b_{2}^{+}(\widehat{Z})$ right-moving and $b^{-}_{2}(\widehat{Z})$ left-moving compact scalar zero modes. Since by assumption $\widehat{Z}$ is a K3-surface we hence end up with $b_{2}^{-}(\widehat{Z})=19$ left-moving scalar modes and $b_{2}^{+}(\widehat{Z})=3$ right-moving scalar modes. These latter scalar zero modes are accompanied by fermionic zero modes and form appropriate $\mathcal{N}=(0,4)$ supermultiplets.
@@
where $\widehat{\phi}^{\mu}$ accounts for all scalar fields in the theory. As in Section 4.2, the interactions are encoded in the field-dependence of $g_{\mu\nu}$. The metric for the scalar fields $\widehat{\mathbf{b}}^{1}$ and $\widehat{u}$ is constant, so that these fields can be treated as free modes. However, due to the nontrivial fibration, the field space metric for the modes arising from $\widehat{B}_{2}$ can have a nontrivial dependence on $\widehat{\mathbf{z}}$ implying that these fields are interacting.
Let us fix a fibre $\widehat{Z}_{0}$ over a point $\widehat{\mathbf{z}}_{0}\in\widehat{\mathbb{P}}^{1}_{b}$. At this point we can consider the scalar fields on the worldsheet obtained by expanding $\widehat{B}_{2}$ in a basis of $H^{2}(\widehat{Z}_{0})$ as
@@
If $\widehat{Z}$ is non-trivially fibred over ^$\widehat{\mathbb{P}}^{1}_{b}$ the basis of 2-forms chosen on $\widehat{Z}_{0}$ undergoes non-trivial monodromies around points in ^$\widehat{\mathbb{P}}^{1}_{b}$. Thus, as we change $\widehat{\mathbf{z}}_{0}$, the basis of 2-forms in which we expand $\widehat{B}_{2}$ changes. As a consequence, the kinetic term for those $\widehat{b}_{2}^{i}$ whose associated $\widehat{\omega}_{i}$ does not correspond to a harmonic 2-form globally defined on $\widehat{V}$ depends on $\widehat{\mathbf{z}}_{0}$. In contrast, if a 2-form in $\widehat{\omega}_{a}\in H^{2}(\widehat{Z})$ is globally defined on $\widehat{V}$, the kinetic term for the corresponding $\widehat{b}^{a}_{2}$ is independent of $\widehat{\mathbf{z}}_{0}$ and in fact completely field-independent. The $\widehat{b}_{2}^{a}$ are hence free fields on the worldsheet. In case $\widehat{\omega}_{a}$ is globally defined on $\widehat{V}$ it is the pull-back of a 2-form in $H^{1,1}(\widehat{V})$ under the inclusion map $\widehat{\iota}:\widehat{Z}\hookrightarrow\widehat{V}$. The lattice $\widehat{\iota}^{*}(H^{1,1}(\widehat{V}))$ is known as the polarization lattice, $\widehat{\Lambda}_{\rm pol}$, of the K3-fibre $\widehat{Z}$ with signature [^37] Note that since $\widehat{Z}$ is nontrivially fibred over base ^$\widehat{\mathbb{P}}^{1}_{b}$, the case $\widehat{\rho}=19$ is excluded.
$\text{sgn}(\widehat{\Lambda}_{\rm pol})=(1,\widehat{\rho}),\qquad 0\leq\widehat{\rho}<19\,,\qquad\widehat{\Lambda}_{\rm pol}=\widehat{\iota}^{*}(H^{1,1}(\widehat{V}))\,.$ (5.5)
@@
$\text{sgn}(\widehat{\Lambda}_{\rm trans})=(2,19-\widehat{\rho})\,.$ (5.6)
For $\widehat{\rho}<18$, the transcendental lattice $\widehat{\Lambda}_{\rm trans}$ contains two left-moving interacting fields which form real scalar fields on the worldsheet once combined with the two right-moving interacting fields. We denote the resulting compact real scalar fields by $(\widehat{\mathbf{b}}^{0},\widehat{{\mathbf{b}}}^{3})$. They can be viewed as the coordinates of a torus in the target space of the sigma-model on the string worldsheet that is nontrivially fibred over ^$\widehat{\mathbb{P}}^{1}_{b}$. [^39] If $\widehat{\rho}=18$ we are in similar situation as discussed in Section 4.3.3. The remaining left-moving interacting scalar fields do not give rise to a massless gauge theory in the bulk.
@@
^$\widehat{\mathcal{M}}=T^{2}_{(\widehat{\mathbf{b}}^{1},\widehat{\mathbf{b}}^{2})}\times\left(T^{2}_{(\widehat{\mathbf{b}}^{0},\widehat{{\mathbf{b}}}^{3})}\to\widehat{\mathbb{P}}^{1}_{\widehat{\mathbf{z}}_{0}}\right)\times\mathbb{C}^{*}_{\widehat{u}}\,.$ (5.7)

### v2

@@ 4.3.2 $\text{rk}(\Lambda_{\rm trans})=3$ or $2$ ($\rho=18$ or $19$)
$\mathcal{M}=T^{2}_{U=T}\times\left(T^{2}_{(\mathbf{b}^{0},\mathbf{c}^{0})}\to\mathbb{C}_{|\mathbf{z}_{0}|e^{{\mathbf{b}}^{3}}}\right)\times\mathbb{C}^{*}_{u=\Phi e^{\arg(\mathbf{z}_{0})}}\,$, (4.39)
@@
$\mathcal{M}=T^{2}_{T=U=\tau_{0}}\times\left(T^{2}_{(\mathbf{b}^{0},\mathbf{c}^{0})}\to\mathbb{C}_{|\mathbf{z}_{0}|e^{{\mathbf{b}}^{3}}}\right)\times\mathbb{C}^{*}_{u=\Phi e^{\arg(\mathbf{z}_{0})}}\,$, (4.40)
where the complex structure and Kähler parameter of the torus are fixed to either $\tau_{0}=i$ or $\tau_{0}=(-1)^{1/3}$. For $\text{rk}(\Lambda_{\rm trans})=2$ we hence obtain a $\sigma$-model corresponding to the heterotic string on a local K3 times a torus with Kähler and complex structure fixed to one of the orbifold points in the fundamental domain of $\mathrm{SL}(2,\mathbb{Z})$. The heterotic dual frame of models with $\rho=19$ including non-perturbative sectors arising from NS5-branes are investigated in more detail in [81].
@@ 4.3.3 $\text{rk}(\Lambda_{\rm trans})=21$ ($\rho=0$)
The final case to discuss is $\rho=0$ (i.e. $b=19$), corresponding to $\text{rk}(\Lambda_{\rm trans})=21$ and $\text{rk}(\Lambda_{\rm pol})=1$. The real scalar field $\mathbf{b}^{3}$ defined in (4.35) now has a free left-moving and an interacting right-moving piece. This leads to a left-right asymmetric worldsheet theory for $\mathbf{b}^{3}$, which cannot correspond to the coordinate on a geometric K3 surface. Accordingly, in $\mathcal{M}$ the geometric $\left(T^{2}_{(\mathbf{b}^{0},\mathbf{c}^{0})}\to\mathbb{C}_{|\mathbf{z}_{0}|e^{{\mathbf{b}}^{3}}}\right)$-factor must be replaced by a non-geometric CFT. Since $b_{-}^{3}$ is now a free field, this CFT realises an additional left-moving ${\rm U}(1)$ current algebra on the worldsheet. In the bulk, the total gauge group is therefore a maximal rank subgroup of $E_{8}\times E_{8}\times\mathrm{U}(1)^{5}$ (or of $\mathrm{Spin}(32)/\mathbb{Z}_{2}\times\mathrm{U}(1)^{5}$). The extra ${\rm U}(1)$ factor suggests that the dynamics is captured by a CFT whose target is at an orbifold point.
@@
The K3 surface that appears in the target space of the $\sigma$-model on the string worldsheet is non-compact since the base of its elliptic fibration is the complex plane $\mathbb{C}_{|\mathbf{z}_{0}|e^{{\bf b}_{3}}}$. On the dual heterotic side one might have expected insteabf bd a compactification on a compact K3 that is an elliptic fibration over $\mathbb{P}^{1}$ as opposed to the complex plane $\mathbb{C}_{|\mathbf{z}_{0}|e^{{\bf b}_{3}}}$. However, our analysis applies to the string at a point in the moduli space corresponding to a type II limit realised as a Tyurin degeneration. This maps to the tensionless limit for the string in Planck units, cf. (3.29). In the dual heterotic frame, the tensionless limit amounts to taking
@@
Thus in Planck units the volume of all curves on K3het diverges in the tensionless limit for the heterotic string, including in particular the base $\mathbb{P}^{1}$ of the elliptically fibred K3het. This explains why the analysis of the $\sigma$-model in the Tyurin degeneration only detects a non-compact base of the target space K3het. By the same reasoning, the volume of the torus fibre of K3het in Planck units must diverge in the emergent string limit. This perfectly matches the divergence of the moduli space metric, (4.19), for the interacting zero modes of the $\sigma$-model, including the zero modes $(\mathbf{b}^{0},\mathbf{c}^{0})$ associated with the elliptic fibre.
@@ 4.3.5 BPS invariants and modularity
[…]
@@
[…]
@@
[…]
@@
where $C_{0}=\partial_{*}\iota_{*}\Gamma_{0}\in\Lambda_{\rm trans}$ denotes the curve on $Z$ over which the asymptotic form of $\Gamma_{0}$ is fibered close to the Tyurin degeneration (cf. the discussion below (3.8)). If ${\rm Gr}_{3}=\emptyset$, $\theta(q)$ is in fact modular.
@@ 5.1 Worldsheet Theory of NS5-Brane String in Type IIA
$\widehat{Z}\ \mathrel{{\mathop{\hookrightarrow}\limits^{\widehat{\iota}}}}$ $\ \ \widehat{V}$ (5.1)
$\ \ \downarrow$
$\ \ \widehat{\mathbb{P}}^{1}_{b}\,.$
The limit in the Kähler moduli space in which $\widehat{\mathbb{P}}^{1}_{b}$ becomes large while the volume of the generic fibre remains constant corresponds to an emergent string limit. The emergent string in this case is given by the NS5-brane wrapping the generic K3-fibre and extended along two of the non-compact directions [8]. The situation is depicted in Figure 4.
Figure 4: A geometric depiction of the set-up described around equation (5.1). The Type IIA NS5-brane (indicated in green) is located at a generic point in the base $\widehat{\mathbb{P}}^{1}_{b}$ and wraps the generic K3-fibre $\widehat{Z}$. The emergent string is obtained in the limit where the volume of the base becomes large, while the volume of the generic fibre stays constant.
@@
[…]
@@
where $\widehat{\phi}^{\mu}$ accounts for all scalar fields in the theory. As in Section 4.2, the interactions are encoded in the field-dependence of $g_{\mu\nu}$. The metric for the scalar fields $\widehat{\mathbf{b}}^{1}$ and $\widehat{u}$ is constant, so that these fields can be treated as free modes. However, due to the nontrivial fibration, the field space metric for the modes arising from $\widehat{B}_{2}$ can have a nontrivial dependence on $\widehat{\mathbf{v}}$ implying that these fields are interacting.
Let us fix a fibre $\widehat{Z}_{0}$ over a point $\widehat{\mathbf{v}}\in\widehat{\mathbb{P}}^{1}_{b}$. At this point we can consider the scalar fields on the worldsheet obtained by expanding $\widehat{B}_{2}$ in a basis of $H^{2}(\widehat{Z}_{0})$ as
@@
[…]
$).\text{sgn}(\widehat{\Lambda}_{\rm pol})=(1,\widehat{\rho}),\qquad 0\leq\widehat{\rho}<19\,,\qquad\widehat{\Lambda}_{\rm pol}=\widehat{\iota}^{*}(H^{1,1}(\widehat{V}$))\,. (5.5)
@@
For $1\leq\widehat{\rho}$, the free right-moving scalar together with one of the free left-moving scalars forms a free compact real scalar on the worldsheet of the string. We denote this real scalar field by $\widehat{\mathbf{b}}^{2}$. It realises a global left-moving and right-moving ${\rm U}(1)$ algebra on the string, leading to a massless $U(1)^{2}$ gauge theory in the bulk associated with an $S^{1}$ target space. Together with the scalar field $\widehat{\mathbf{b}}^{1}$ this constitutes a target space factor $T^{2}_{(\widehat{\mathbf{b}}^{1},\widehat{\mathbf{b}}^{2})}$. The other free compact left-moving modes realise a rank $\widehat{\rho}-1$ current algebra on the string worldsheet, leading to a gauge theory in the bulk with gauge group of rank $\widehat{\rho}-1$}. [^39] If $\widehat{\rho}=0$, the $S^{1}$ radius is fixed to the self-dual radius and the left-moving ${\rm U}(1)$ is broken similar to our discussion in Section 4.3.2.
@@
For $\widehat{\rho}<18$, the transcendental lattice $\widehat{\Lambda}_{\rm trans}$ contains two left-moving interacting fields which form real scalar fields on the worldsheet once combined with the two right-moving interacting fields. We denote the resulting compact real scalar fields by $(\widehat{\mathbf{b}}^{0},\widehat{{\mathbf{c}}}^{0})$. They can be viewed as the coordinates of a torus in the target space of the sigma-model on the string worldsheet that is nontrivially fibred over $\widehat{\mathbb{P}}^{1}_{b}$. [^40] If $\widehat{\rho}=18$ we are in similar situation as discussed in Section 4.3.3. The remaining left-moving interacting scalar fields do not give rise to a massless gauge theory in the bulk.
@@
$\widehat{\mathcal{M}}=T^{2}_{(\widehat{\mathbf{b}}^{1},\widehat{\mathbf{b}}^{2})}\times\left(T^{2}_{(\widehat{\mathbf{b}}^{0},\widehat{\mathbf{c}}^{0})}\to\widehat{\mathbb{P}}^{1}_{\hat{\mathbf{v}}}\right)\times\mathbb{C}^{*}_{\widehat{u}}\,$. (5.7)